# OPUS: Occupancy Prediction Using a Sparse Set

### v1

@@ 1 Introduction
Compared with well-established box representations [7, 22, 19, 35, 28, 44], voxel based occupancy [15, 33, 9, 34] can provide finer grinded geometry and semantic information for the surrounding scene. For example, it is not straightforward to use bounding boxes to describe vehicles with doors open or cranes with outriggers deployed. While occupancy can naturally describe such uncommon shapes. Thus occupancy prediction is quickly gaining traction in the autonomous driving community.
Recent approaches [3, 42, 8, 26, 15] to the task predominantly rely on dense data representation, with a direct one-to-one correspondence between feature points and physical voxels. It has come to our attention that the vast majority of physical voxels is empty. For instance, in SemanticKITTI [1], approximately 67% of all voxels are empty, while in Occ3D-nuScenes [34], this proportion exceeds 90%. Such sparse nature of occupancy data renders the direct dense representation undeniably inefficient, as majority of the computation is allocated towards empty voxels. Alternative sparse latent representations have been explored to alleviate such inefficiency, such as the Tri-Perspective View representation [33, 8] or reduced solution spaces [20, 9], leading to notably reduced computational costs. However, these approaches still treat occupancy prediction as a classification problem at specific locations, necessitating complex intermediate designs and explicit modeling of 3D spaces.
@@ 2.1 Occupancy prediction
Occupancy prediction entails determining the occupancy status and class of each voxel within a 3D space. This task has recently become a foundational perception task in autonomous driving and raises great interests from both academic and industrial communities. Conventional methods [3, 42, 8, 26, 15, 37, 34] typically employ the continuous and dense feature representation, which, however, suffer from computational redundancy due to the inherent sparsity of occupancy data. In addressing this issue, Tang et al. [33] compresses the dense feature using the Tri-Perspective View representation for model efficiency. Recently, several transformer-based approaches [20, 9, 13] with sparse queries have emerged. For example, OccupancyDETR [9] conducts object detection followed by assigning each object with one query for occupancy completion. VoxFormer [13] generates 3D voxels from a set of sparse queries, corresponding to occupied locations identified through a pre-task of depth estimation. Meanwhile, SparseOcc [20] employs a series of sparse voxel decoders to filter out empty grids and predict occupied statuses of retained voxels in each stage. While these approaches have succeeded in reducing computational costs, they often necessitate multi-stage processes and intricate space modeling. In contrast, our method directly applies sparse queries to regress the occupied locations without pre-defined locations, facilitating an elegant and end-to-end occupancy prediction.
@@ D.1 Comparison of Hungarian matching and our method
D.3 Detailed structure of OPUS decoder
Table 6: Comparisons between different sparsification strategies.
SparseOcc (4000/16000/64000) 28.4 34.9 39.6 34.3 17.3
PETR v2 2500 256 24.4 31.0 36.3 30.6 13.8
OPUS 2400 32 31.7 38.8 43.4 38.0 13.4
D.4 Comparisons between different sparsification strategies
In Tab. 6, we compare OPUS to two other models with different sparsification strategies. The first baseline is SparseOcc, which achieves sparsification by filtering out empty voxels at various cascade stages. Following PETRv2 [23], the second baseline is a pillar-patch based method that partitions the 3D space into a small number of pillar-patches. We use $50\times 50$ queries with each corresponding to the classification of neighbouring $4\times 4\times 16$ voxels. The sparsification strategy in OPUS involves two task: classification and regression whereas the other two sparsification methods only include the classification task. For a fair comparison, all these models are trained for 100 epochs. From this table and Tab. 6, we observe that increasing training epochs from 24 to 100 brings minor improvements to SparseOcc, likely due to performance bottleneck caused by cumulative filtering errors at various stages. Utilizing sparsification method like PETR v2 fails to achieve better results, with 30.6 $\text{RayIoU}_{\text{}}$ and 13.8 FPS. In contrast, our model achieve best results after sufficient training with a $\text{RayIoU}_{\text{}}$ score of 38.0,far outperforming SparseOcc with a $\text{RayIoU}_{\text{}}$ score of 34.3. On the other hand, our model can also runs in a real-time speed. These results demonstrates the superiority of our sparification procedure.
@@ D.4 Comparisons between different sparsification strategies
D.5 Effects of various refined points number in last layer.
@@
D.6 Predictions across different distances
@@
Table 9: Performance on the Waymo-Occ3D dataset.
Model
General
Vehicle
Bicyclist
Ped.
Sign
Tfc. light
Pole
Cons. cone
Bicycle
Motorcycle
Building
Vegetaion
Treetrunk
Road
Sidewalk
mIoU
$\text{RayIoU}_{\text{}}$
FPS
D.7 Experiment on the Waymo-Occ3D dataset.
We further simply implement OPUS on the Waymo-Occ3D [31] dataset to explore the generalization and robustness of OPUS. As Waymo-Occ3D is not commonly used as a standard benchmark for vision-centric approaches, the only vision-based method we found with reported results on this dataset is the Occ3D paper, which evaluates BEVDet, TPVFormer, BEVFormer, and the newly proposed CTF-Occ [34]. We trained the OPUS-L (1f) on 20% of the dataset for a fair comparison with these baselines. As reported in Tab. 9, despite not fine-tuning the training configurations, OPUS-L already achieves 19.0 mIoU, outperforming all previous methods. Moreover, OPUS-L also reaches 8.5 FPS on the Waymo-Occ3D dataset, which is around 3 times the speed of CTF-Occ and 2 times the speed of BEVFormer.
@@ E.1 Differences between SparseOcc and OPUS
Figure 7: Illustration of safety threat due to discrepancies between evaluation metrics and real-world scenarios. (a) Before evaluation, the camera visibility mask is first generated according to camera intrinsics and extrinsics, as well as the ground-truth occupancy. Then, the dense prediction will be masked by the camera visibility mask to get the final prediction for evaluation. (b) For real-world usage, we cannot have camera visibility reasoning without knowing the ground-truth occupancy. We can only generate the view mask from camera intrinsics and extrinsics, which fails to filter out the over-estimated voxels from dense models. (c) Plenty of false positive predictions are made close to the ego vehicle, marked by the symbol of red star. These erroneously predicted voxels are filtered during evaluating mIoU, but could cause hazardous safety issue in real-world usage. (d) The OPUS produces sparse occupancy predictions and suffers much less from the over-estimation. Consequently, no such safety threat occurs in this scenario. Best viewed in color.
@@
E.3 Visualization of the self-attention.
In Fig. 9, we visualize the self-attention procedure in decoders. Specifically, after selecting certain queries, we identify the points with the top 10 attention weights and project them onto 2D images for clear visualization. Generally speaking, the self-attention scheme tends to highlight queries representing the same object, while maintaining relative diversity within the instances. For instance, the highlighted points adaptively cover most of the vehicle areas in the last two images, enabling the query to gather more discriminative features. Additionally, from the visualization, we observe that attention weights tend to decrease with distances from the pivot point. This observation is consistent across different cases, reflecting a rational pattern in the attention mechanism as closer points tend to have more relevant feature representations.
E.4 Occupancy predictions of different methods.
In Fig. 10, We further provide more visualizations of occupancy predicted by FB-Occ, SparseOcc, and proposed OPUS.
E.5 Failure cases
As shown in Fig. 3 and Fig. 10, a common OPUS failure mode is the prediction of scattered and discontinuous surfaces at long distances. Another is the presence of holes in predicted driving surface, a phenomenon also observed in SparseOcc due to the sparsity properties.

### v2

@@ 1 Introduction
Compared with well-established box representations [8, 23, 20, 39, 29, 48, 32], voxel based occupancy [16, 37, 10, 38, 33] can provide finer geometry and semantic information for the surrounding scene. For example, it is not straightforward to use bounding boxes to describe vehicles with doors open or cranes with outriggers deployed. While occupancy can naturally describe such uncommon shapes. Thus occupancy prediction is quickly gaining traction in the autonomous driving community.
Recent approaches [3, 46, 9, 27, 16, 34] to the task predominantly rely on dense data representation, with a direct one-to-one correspondence between feature points and physical voxels. It has come to our attention that the vast majority of physical voxels is empty. For instance, in SemanticKITTI [1], approximately 67% of all voxels are empty, while in Occ3D-nuScenes [38], this proportion exceeds 90%. Such sparse nature of occupancy data renders the direct dense representation undeniably inefficient, as majority of the computation is allocated towards empty voxels. Alternative sparse latent representations have been explored to alleviate such inefficiency, such as the Tri-Perspective View representation [37, 9] or reduced solution spaces [21, 10], leading to notably reduced computational costs. However, these approaches still treat occupancy prediction as a classification problem at specific locations, necessitating complex intermediate designs and explicit modeling of 3D spaces.
@@ 2.1 Occupancy prediction
Occupancy prediction entails determining the occupancy status and class of each voxel within a 3D space. This task has recently become a foundational perception task in autonomous driving and raises great interests from both academic and industrial communities. Conventional methods [3, 46, 9, 27, 16, 41, 38, 4] typically employ the continuous and dense feature representation, which, however, suffer from computational redundancy due to the inherent sparsity of occupancy data. In addressing this issue, Tang et al. [37] compresses the dense feature using the Tri-Perspective View representation for model efficiency. Recently, several transformer-based approaches [21, 10, 14] with sparse queries have emerged. For example, OccupancyDETR [10] conducts object detection followed by assigning each object with one query for occupancy completion. VoxFormer [14] generates 3D voxels from a set of sparse queries, corresponding to occupied locations identified through a pre-task of depth estimation. Meanwhile, SparseOcc [21] employs a series of sparse voxel decoders to filter out empty grids and predict occupied statuses of retained voxels in each stage. While these approaches have succeeded in reducing computational costs, they often necessitate multi-stage processes and intricate space modeling. In contrast, our method directly applies sparse queries to regress the occupied locations without pre-defined locations, facilitating an elegant and end-to-end occupancy prediction.
@@ 4.3 Ablation study and visualizations
Visualization of the self-attention. For better understanding what query points attend to, we visualize top 10 query points with highest self attention weights for each query. We project them on 2D image for better visualization. Here are some interesting finding from Fig. 7. Generally speaking, the query tends to attend to neighbouring query points. For example, the sidewalk query in the first image and the car query in the second image, allow local information from neighbouring query points flow into, enabling the query to capture detailed local information. Additionally, the query maintains the ability to attend to semantic related locations even if they are not very close to the query point. For instance, the vegetation query in the first image of Fig. 7 not only attends to the stem of the tree, but also the grass, indicating that the query can capture semantically related information for more accurate predictions. Another notable observation is that query can sense terrain-related information. For instance, in the third image, the sidewalk query points attend to other points along a straight line following the road’s edge, highlighting the model’s ability to understand scene related structure of the environment.
Comparisons between different sparsification strategies In Tab. 4, we compare OPUS to two other models with different sparsification strategies. The first baseline is SparseOcc, which achieves sparsification by filtering out empty voxels at various cascade stages. Following PETRv2 [24], the second baseline is a pillar-patch based method that partitions the 3D space into a small number of pillar-patches. We use $50\times 50$ queries with each corresponding to the classification of neighbouring $4\times 4\times 16$ voxels. For a fair comparison, all these models are trained for 100 epochs. In contrast, our model achieve best results after sufficient training with a $\text{RayIoU}_{\text{}}$ score of 38.0,far outperforming SparseOcc with a $\text{RayIoU}_{\text{}}$ score of 34.3. On the other hand, our model can also runs in a real-time speed. These results demonstrates the superiority of our sparification procedure.
Comparisons on the Waymo-Occ3D dataset. We further simply implement OPUS on the Waymo-Occ3D [35] dataset to explore the generalization and robustness of OPUS. As Waymo-Occ3D is not commonly used as a standard benchmark for vision-centric approaches, the only vision-based method we found with reported results on this dataset is the Occ3D paper, which evaluates BEVDet, TPVFormer, BEVFormer, and the newly proposed CTF-Occ [38]. We trained the OPUS-L (1f) on 20% of the dataset for a fair comparison with these baselines. As reported in Tab. 5, despite not fine-tuning the training configurations, OPUS-L already achieves 19.0 mIoU, outperforming all previous methods. Moreover, OPUS-L also reaches 8.5 FPS on the Waymo-Occ3D dataset, which is around 3 times the speed of CTF-Occ and 2 times the speed of BEVFormer.
Table 4: Comparisons between different sparsification strategies.
SparseOcc (4000/16000/64000) 28.4 34.9 39.6 34.3 17.3
PETR v2 2500 256 24.4 31.0 36.3 30.6 13.8
OPUS 2400 32 31.7 38.8 43.4 38.0 13.4
Table 5: Performance on the Waymo-Occ3D dataset.
Model
General
Vehicle
Bicyclist
Ped.
Sign
Tfc. light
Pole
Cons. cone
Bicycle
Motorcycle
Building
Vegetaion
Treetrunk
Road
Sidewalk
mIoU
$\text{RayIoU}_{\text{}}$
FPS
@@ 5 Conclusions and limitations
Acknowledgments
This research was supported by NSFC (NO. 62225604, NO. 62276145), the Fundamental Research Funds for the Central Universities (Nankai University, 070-63223049) and partially supported by NSFC (No. 62376153, 62402318, 24Z990200676). Computations were supported by the Supercomputing Center of Nankai University (NKSC).
@@ D.1 Comparison of Hungarian matching and our method
D.3 Effects of various refined points number in last layer.
@@
D.4 Predictions across different distances
@@ E.1 Differences between SparseOcc and OPUS
Figure 8: Illustration of safety threat due to discrepancies between evaluation metrics and real-world scenarios. (a) Before evaluation, the camera visibility mask is first generated according to camera intrinsics and extrinsics. Then, the dense prediction will be masked to get the final prediction for evaluation. (b) For real-world usage, we cannot have camera visibility reasoning without knowing the ground-truth occupancy. We can only generate the view mask from camera intrinsics and extrinsics, which fails to filter out the over-estimated voxels from dense models. (c) Plenty of false positive predictions are made close to the ego vehicle, marked by the symbol of red star. These erroneously predicted voxels are filtered during evaluating mIoU, but could cause hazardous safety issue. (d) The OPUS produces sparse occupancy predictions and suffers much less from the over-estimation. Consequently, no such safety threat occurs in this scenario. Best viewed in color.
@@
E.3 Occupancy predictions of different methods.
In Fig. 10, We further provide more visualizations of occupancy predicted by FB-Occ, SparseOcc, and proposed OPUS. As shown in Fig. 3 and Fig. 10, a common OPUS failure mode is the prediction of scattered and discontinuous surfaces at long distances. Another is the presence of holes in predicted driving surface, a phenomenon also observed in SparseOcc due to the sparsity properties.